show clear working Konrad Company reported the following operating results: Sales Variable Costs Contribution Margin Fixed Costs Operating IncomHang Ten produces sport socks. The company has fixed expenses of $90,000 and variable expenses of $0.90 per package. Each pac
Answer #1

Part 1

First calculate operating leverage factor

Operating leverage factor = contribution margin / operating income = 128000/40000 = 3.2

Increase in operating income = increase in sales volume * operating leverage factor = 12%*3.2 = 38.4%

(However, this option is not given, 41.6% is answer if the increase in sales is 13%, but as nearest to the correct answer is 41.6%. it is selected or if possible this should be reported)

Part 2

Desired sales = (fixed expenses + operating income)/contribution margin per unit

Contribution margin per unit = sales price per unit – variable costs per unit

Desired sales = (105000+29000)/(1.80-0.70) = 121819 units

Would have to sell = 121819-132223 = 10404 units fewer.